# Universidade de Lisboa, Av. Rovisco Pais 1, 1049-001 Lisboa, Portugal

### v3

@@ 9 Sampling Methods
At this stage, we do not impose the bounds from Dark Matter direct and indirect searches. Such bounds rule out most of the few points we found with acceptable relic density, with Fig.4 showing the comparison of the points sampled near the alignment limit with the most recent exclusion line from the LZ collaboration LZ:2024zvo. As most points sampled do not have a significant relic density, we continue with a new Machine Learning method instead of continuing with longer sets of inefficient traditional sampling.
@@
Figure 4: Direct detection results for the scan near the alignment limit $\alpha_{1}=-[0.9,1.1]\beta$, with the relevant quantity for nucleon scattering $\sigma_{SI}^{Xe}$ obtained from eq. (132). The color code is the same as Fig. 2, with the addition of the LZ exclusion line from 2024 LZ:2024zvo drawn in blue.
We followed the procedure for efficiently sampling the parameter space shown in deSouza:2022uhk and further developed in Romao:2024gjx; deSouza:2025uxb; deSouza:2025bpl  [^5] Alternative approaches to successfully sample the parameter space of BSM models have been developed in cases where training datasets are available Caron:2019xkx; Hollingsworth:2021sii; Goodsell:2022beo; Diaz:2024yfu; Batra:2025amk; Hammad:2025wst. For more broad applications of Machine Learning in the field of particle physics, consider the comprehensive reviews Feickert:2021ajf; Plehn:2022ftl.. The first step is to define constraint functions, $C(\mathcal{O})$, as
@@
The optimization algorithm is the Covariant Matrix Adaptation Evolutionary Strategy (CMA-ES) Hansen2001; hansen2023cmaevolutionstrategytutorial, characterized by an iterative sampling according to a multivariate normal distribution, initialized with its mean at a random point in parameter space and its covariance matrix set to the identity matrix, $\mathbb{1}$, scaled by a constant. A generation of candidate solutions is sampled from this distribution and the candidates are ranked from best to worst based on the loss function in eq. (136). The best candidates are then used to compute a new mean and approximate the covariance matrix, for the next iterative step.
CMAES has one critical downside: it has a limited exploration capacity due to the highly localized nature of the algorithm. To mitigate this, similarly to Ref. Romao:2024gjx, we implemented a novelty reward into the loss function. This is done by adding the density of found points as a penalty to the loss function, thus allowing it to be minimal when the point is good and away from other good points. To this end, we used the Histogram Based Outlier System (HBOS), which computes a density penalty between 0 and 1, such that a new point has 0 and a point similar to previous ones approach the maximum 1, basing itself on the abundance values for chosen parameters.
@@
Figure 5: Results from a run using the optimization algorithm CMA-ES, without the novelty reward or seeded runs methods described in the text. In the Left panel we show the rapid convergence towards the imposed interval on the total relic density, $\Omega_{T}h^{2}\in[0.1164,0.1236]$, and in the Right panel the obtained value for the direct detection quantity $\sigma_{\text{SI}}^{\text{Xe}}\xi$, with the most recent constraint from LZ LZ:2024zvo, shown as the blue line. The points in red fail one of the constraints not related to DM and the points in green meet all constraints except DM. The points with a blue color are points in green that also satisfy the correct relic density and direct detection scattering bounds.
@@ 10 Final Machine Learning Results
When we include indirect constraints, with exclusion lines shown in Fig. 1, we observe an increase in the time per generation of a factor of 4 and an increased convergence time. The strategy followed was then to only use micrOMEGAs-6.2.3 Alguero:2023zol to numerically calculate the relic density and direct detection variables. For a set of simulations, we took parameter points that met the relic density from Planck and nucleon scattering constraints to check the annihilation bounds. As a next step, we passed again the valid (relic and direct detection) points through micrOMEGAs-6.2.3, probing for consistency with indirect detection bounds. We found the large majority of points to already satisfy the current indirect detection constraints. The final results, also satisfying all other constraints described in Sections 48, have the Dark Matter relic density within $3\sigma$ of the limits obtained by the Planck experiment Planck:2018vyg, satisfy the most recent direct detection bounds from LZ LZ:2024zvo, as well as the indirect detection constraints from Fig. 1. All other constraints described in Sections 48 are also considered.
@@
No valid points were found in the mass region below $m_{h_{1}}/2$, as the latest LHC bound on the invisible branching ratio of the Higgs, BR($h\rightarrow\text{inv}$) $\leq$ 0.107 at $95\%$ C.L. ATLAS:2023tkt, comes into effect.
In Fig. 6, we show the constraints from direct detection [^6] The data files and general layout for the plots was obtained from the public repository by Ciaran O’Hare avaliable in https://github.com/cajohare/DirectDetectionPlots. for the mass region studied. The solid lines shown correspond to the most recent exclusion bounds from experiments, with XENONnT XENON:2025vwd, PandaX-4T PandaX:2024qfu and LZ LZ:2024zvo. The dashed lines show the projections for DarkSide-20k DarkSide-20k:2017zyg and XLZD XLZD:2024nsu. The neutrino floor is shown in grey, as defined in Ref. OHare:2021utq.
Figure 6: The red region correponds to valid points generated with our model. The solid lines shown correspond to the most recent exclusion bounds from experiments, with XENONnT XENON:2025vwd, PandaX-4T PandaX:2024qfu and LZ LZ:2024zvo. The dashed lines show the projections for DarkSide-20k DarkSide-20k:2017zyg and XLZD XLZD:2024nsu. The neutrino floor is shown in grey, as defined in Ref. OHare:2021utq.
As one can see from Fig. 6, there are many valid points obeying the current LZ bounds (in addition to matching the correct relic density). Some of those points could be excluded (or confirmed) by future experiments, such as DarkSide-20k or XLZD. However, the fact that many red (valid model) points lie within the grey (neutrino fog) region, implies that this model will be hard to exclude with future direct detection experiments  [^7] There is also the possibility of improving theoretical and experimental tools in order to break the degeneracy between neutrino backgrounds and potential WIMP signals Davis:2014ama; Gelmini:2018ogy; Arcadi:2025ifl; Vahsen:2020pzb..
@@ 11 Conclusion
Points in parameter space where searched for, which obeyed these conditions and, in addition, satisfy current experimental bounds arising from the oblique radiative parameters, current LHC bounds on the $125\textrm{GeV}$ Higgs as well as searches for additional scalar particles, and also direct, indirect, and relic abundance constraints. These constraints have been implemented using HiggsTools-1.1.3 Bahl:2022igd and micrOMEGAs-6.2.3 Alguero:2023zol.
Three different scanning strategies have been utilized: i)  a random scan without any prior assumptions on the parameter space; ii)  a scan close to the alignment limit of the 2HDM, defined as $\alpha_{1}=-\beta$ for the choices in eqs. (13) and (15); iii)  employing the Artificial Intelligence black box optimization approach. We have found that the random scan and the scan close to the alignment limit of the 2HDM produce similar results, although (of course) the latter is more efficient than the first, by a factor close to three. Neither of these produce points with the correct relic density in a reasonable time frame. Thus, we turned to a machine learning method. Using unseeded runs, we quickly find solutions with reasonable relic density, but yielding limited exploration of the various mass regions. In contrast, using a novelty reward strategy and seeded runs, we are able to find valid points in a larger whole mass region: $125\textrm{GeV}/2<m_{DM}\lesssim 800\textrm{GeV}$.

### v4

@@ 9 Sampling Methods
At this stage, we do not impose the bounds from Dark Matter direct and indirect searches. Such bounds rule out most of the few points we found with acceptable relic density, with Fig.4 showing the comparison of the points sampled near the alignment limit with the most recent exclusion line from the LZ collaboration Aalbers and others (2024a). As most points sampled do not have a significant relic density, we continue with a new Machine Learning method instead of continuing with longer sets of inefficient traditional sampling.
@@
Figure 4: Direct detection results for the scan near the alignment limit $\alpha_{1}=-[0.9,1.1]\beta$, with the relevant quantity for nucleon scattering $\sigma_{SI}^{Xe}$ obtained from eq. (132). The color code is the same as Fig. 2, with the addition of the LZ exclusion line from 2024 Aalbers and others (2024a) drawn in blue.
We followed the procedure for efficiently sampling the parameter space shown in de Souza et al. (2023) and further developed in Romão and Crispim Romão (2024); de Souza et al. (2025b, a)  [^5] Alternative approaches to successfully sample the parameter space of BSM models have been developed in cases where training datasets are available Caron et al. (2019); Hollingsworth et al. (2021); Goodsell and Joury (2023); Diaz et al. (2024); Batra et al. (2025); Hammad et al. (2025). For more broad applications of Machine Learning in the field of particle physics, consider the comprehensive reviews Feickert and Nachman (2021); Plehn et al. (2022).. The first step is to define constraint functions, $C(\mathcal{O})$, as
@@
The optimization algorithm is the Covariant Matrix Adaptation Evolutionary Strategy (CMA-ES) Hansen and Ostermeier (2001); Hansen (2023), characterized by an iterative sampling according to a multivariate normal distribution, initialized with its mean at a random point in parameter space and its covariance matrix set to the identity matrix, $\mathbb{1}$, scaled by a constant. A generation of candidate solutions is sampled from this distribution and the candidates are ranked from best to worst based on the loss function in eq. (136). The best candidates are then used to compute a new mean and approximate the covariance matrix, for the next iterative step.
CMAES has one critical downside: it has a limited exploration capacity due to the highly localized nature of the algorithm. To mitigate this, similarly to Ref. Romão and Crispim Romão (2024), we implemented a novelty reward into the loss function. This is done by adding the density of found points as a penalty to the loss function, thus allowing it to be minimal when the point is good and away from other good points. To this end, we used the Histogram Based Outlier System (HBOS), which computes a density penalty between 0 and 1, such that a new point has 0 and a point similar to previous ones approach the maximum 1, basing itself on the abundance values for chosen parameters.
@@
Figure 5: Results from a run using the optimization algorithm CMA-ES, without the novelty reward or seeded runs methods described in the text. In the Left panel we show the rapid convergence towards the imposed interval on the total relic density, $\Omega_{T}h^{2}\in[0.1164,0.1236]$, and in the Right panel the obtained value for the direct detection quantity $\sigma_{\text{SI}}^{\text{Xe}}$ given by eq. (132), with the most recent constraint from LZ Aalbers and others (2024a), shown as the blue line. The points in red fail one of the constraints not related to DM and the points in green meet all constraints except DM. The points with a blue color are points in green that also satisfy the correct relic density and direct detection scattering bounds.
@@ 10 Final Machine Learning Results
When we include indirect constraints, with exclusion lines shown in Fig. 1, we observe an increase in the time per generation of a factor of 4 and an increased convergence time. The strategy followed was then to only use micrOMEGAs-6.2.3 Alguero et al. (2024) to numerically calculate the relic density and direct detection variables. For a set of simulations, we took parameter points that met the relic density from Planck and nucleon scattering constraints to check the annihilation bounds. As a next step, we passed again the valid (relic and direct detection) points through micrOMEGAs-6.2.3, probing for consistency with indirect detection bounds. We found the large majority of points to already satisfy the current indirect detection constraints. The final results, also satisfying all other constraints described in Sections 48, have the Dark Matter relic density within $3\sigma$ of the limits obtained by the Planck experiment Aghanim and others (2020), satisfy the most recent direct detection bounds from LZ Aalbers and others (2024a), as well as the indirect detection constraints from Fig. 1. All other constraints described in Sections 48 are also considered.
@@
No valid points were found in the mass region below $m_{h_{1}}/2$, as the latest LHC bound on the invisible branching ratio of the Higgs, BR($h\rightarrow\text{inv}$) $\leq$ 0.107 at $95\%$ C.L. Aad and others (2024), comes into effect.
In Fig. 6, we show the constraints from direct detection [^6] The data files and general layout for the plots was obtained from the public repository by Ciaran O’Hare avaliable in https://github.com/cajohare/DirectDetectionPlots. for the mass region studied. The solid lines shown correspond to the most recent exclusion bounds from experiments, with XENONnT Aprile and others (2025), PandaX-4T Bo and others (2025) and LZ Aalbers and others (2024a). The dashed lines show the projections for DarkSide-20k Aalseth and others (2018) and XLZD Aalbers and others (2024b). The neutrino floor is shown in grey, as defined in Ref. O’Hare (2021).
Figure 6: The red region correponds to valid points generated with our model. The solid lines shown correspond to the most recent exclusion bounds from experiments, with XENONnT Aprile and others (2025), PandaX-4T Bo and others (2025) and LZ Aalbers and others (2024a). The dashed lines show the projections for DarkSide-20k Aalseth and others (2018) and XLZD Aalbers and others (2024b). The neutrino floor is shown in grey, as defined in Ref. O’Hare (2021).
As one can see from Fig. 6, there are many valid points obeying the current LZ bounds (in addition to matching the correct relic density). Some of those points could be excluded (or confirmed) by future experiments, such as DarkSide-20k or XLZD. However, the fact that many red (valid model) points lie within the grey (neutrino fog) region, implies that this model will be hard to exclude with future direct detection experiments  [^7] There is also the possibility of improving theoretical and experimental tools in order to break the degeneracy between neutrino backgrounds and potential WIMP signals Davis (2015); Gelmini et al. (2018); Arcadi et al. (2025); Vahsen and others (2020)..
@@ 11 Conclusion
Points in parameter space where searched for, which obeyed these conditions and, in addition, satisfy current experimental bounds arising from the oblique radiative parameters, current LHC bounds on the $125\textrm{GeV}$ Higgs as well as searches for additional scalar particles, and also direct, indirect, and relic abundance constraints. These constraints have been implemented using HiggsTools-1.1.3 Bahl et al. (2023) and micrOMEGAs-6.2.3 Alguero et al. (2024).
Three different scanning strategies have been utilized: i)  a random scan without any prior assumptions on the parameter space; ii)  a scan close to the alignment limit of the 2HDM, defined as $\alpha_{1}=-\beta$ for the choices in eqs. (13) and (15); iii)  employing the Artificial Intelligence black box optimization approach. We have found that the random scan and the scan close to the alignment limit of the 2HDM produce similar results, although (of course) the latter is more efficient than the first, by a factor close to three. Neither of these produce points with the correct relic density in a reasonable time frame. Thus, we turned to a machine learning method. Using unseeded runs, we quickly find solutions with reasonable relic density, but yielding limited exploration of the various mass regions. In contrast, using a novelty reward strategy and seeded runs, we are able to find valid points in a larger whole mass region: $125\textrm{GeV}/2<m_{DM}\lesssim 1000\textrm{GeV}$.